OpenStudy (anonymous):

graph n=6 in the geometric sequence 40, 20, 10, 5 and see if the values will ever become zero or negative?

OpenStudy (amistre64):

looks like we need to know the general setup for a geo progression. you know of any?

OpenStudy (anonymous):

what do you mean

OpenStudy (amistre64):

i mean, there is a general equation that helps us find values for the sequence.

OpenStudy (anonymous):

ya it wants me to make that equation

OpenStudy (amistre64):

how do we make a geometric sequence? is it something we add to each term? multiply to each term? or something else ....

OpenStudy (anonymous):

since its a geometric sequence you multiply somthing to each term to get the value

OpenStudy (amistre64):

good, lets start there we know the 1st term is 40 40*r = 20, what is the value of r?

OpenStudy (anonymous):

it can be division to so it 40 divided by 2 is 20 and 20 divided by 2 is 10

OpenStudy (amistre64):

so r = 1/2 seems fine now we have a pattern that develops 40r = 20 40rr = 10 40rrr = 5 40r^(n-1) = nth term

OpenStudy (amistre64):

ploting the points on a graph i assume is simple enough to do right? its the 0 or negative part that might be troublesome

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

what rules or methods do you have to accomplish this with. limit ratio? or some other contraption?

OpenStudy (anonymous):

you have to use the equation itself to be able to plot the points since its a geometric you multiplied or divided something to get the values which was 2

OpenStudy (amistre64):

right, we plot 1,40 2,20 3,10 4,5 5,2.5 6,1.25 logically, if we have something, and take half of it away .... we still have something left over correct?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

then since we will always have something left over, then we never get to zero, or less than 0.

OpenStudy (anonymous):

ya but why